Weight of carbon di oxide prepeared by heating 200kg of limestone which is 95% pure

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
CaCO3 - >CaO+Co2 
atomic wt. of CaCO3=40+12+16*3 
=100g 
molec.wt of CaO=40+16=56g 
amt of limestone available=95/100*200=190g 
at the point when CaCO3 is 100g thenCaO is 56g 
at the point when limestone is 190g then CaO is 56/100 *190=106.40g.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)Cheers!!
 
Regards,
Vikas (B. Tech. 4th year
Thapar University)